OpenStudy (anonymous):

Plutonium-240 decays according to the function Q(t)=Q(subscripto)e^(-kt) where Q represents the quantity remaining after t years and k is the decay constant, 0.00011... How long will it take 27 grams of plutonium-240 to decay to 9 grams?

OpenStudy (anonymous):

A. 18,900 years B. 2,100 years C. 1.44 years D. 9,990 years

OpenStudy (campbell_st):

you have\[Q(t) 9====>Q _{0}= 27 =====>k = 0.00011\] then \[9 = 27 e^-0.00011t\] or \[\frac{1}{3} = e^{-0.00011t}\] take the base e log of both sides \[\ln(\frac{1}{3}) = -0.00011t\] \[t = \frac{\ln(\frac{1}{3})}{-0.00011}\] all you need to do is evaluate t